Can someone please the time to help me solve this problem: In a right triangle, the length of the hypotenuse is 5 and the length of one of the legs is 3, determine the tangants of all three sides. I would be really appreciate it if you did help me.
Thank you

Answer by stanbon(75887)

Can someone please the time to help me solve this problem: In a right triangle, the length of the hypotenuse is 5 and the length of one of the legs is 3, determine the tangants of all three sides. I would be really appreciate it if you did help me.
------------------------
Draw the triangle and label the hypotenuse as 5 and one of the legs as 3
Let the 3rd leg be "x".
Then x^2 + 3^2 = 5^2
x^2 = 5^2 - 3^2
x^2 = 25-9
x^2 = 16
x = 4
-----------
Label the "x"-side with a 4.
--------------------
The "tangents of all three sides" is a rather vague qustion.
You usually find the tangent of an angle, not a side.
Maybe you could clarify what you mean by the question.
=================
Cheers,
Stan H.